OpenStudy (igreen):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 2 inches, and the height of the cone is 3 inches. If the diameter of the bubble gum is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? Use 3.14 for pi.

OpenStudy (igreen):

So first I’m going to find the area of the cone, then subtract the area of the bubble gum from it. Volume of the cone will be \(\sf \dfrac{1}{3} \pi r^2 h \rightarrow (\dfrac{1}{3})(3.14)(2^2)(3) \approx 12.56\). Volume of Bubble Gum will be \(\sf \dfrac{4}{3} \pi r^3 \rightarrow (\dfrac{4}{3})(3.14)(0.25^3) \approx 0.065\) So therefore we can have \(\sf 12.56 - 0.065\) or approximately \(\sf 12.50\) cubic inches for flavored ice.

OpenStudy (igreen):

I kinda wanna be careful..because there's also an option labeled \(\sf 12.60\)..

OpenStudy (igreen):

@Directrix

OpenStudy (igreen):

Yeah I'm pretty sure \(\sf 12.50\) is correct.

OpenStudy (anonymous):

@hartnn :p

Directrix (directrix):

I think I and the Asker worked this same problem a year ago. Take a look at our work: http://openstudy.com/study#/updates/5285a807e4b0b3376510ce26 @iGreen

Directrix (directrix):

The numbers are different - not the same problem - but maybe the strategy is the same. I remember wondering if the volume at the tip of the cone should be included.

OpenStudy (igreen):

I'll just go with my answer..thanks anyway.